The me-tools Package

R W Kaszeta

July 10, 1998

Introduction

The me-tools package provides a number of macro definitions useful for mechanical engineers.

Currently, the package provides macros for the following:

1. Convenient typesetting of non-dimensional numbers such as Re and Pr.

2. Macros to streamline the typsetting of many fraction-like items like total and partial derivatives.

1 Non-dimensional Numbers

The me-tools package provides a number of macros for typesetting non-dimensional numbers, listed in table 1

All of these macros are defined so that they work both inside and outside of math mode (i.e. \RE and $\RE$ both produce Re) so you can refer to these numbers in a sentence without needing to enter math mode.

Additionally, the user can create their own macros using the \newnondim command, with the first argument being the macro name and the second the (math-mode) definition for the label, without formatting. For example, to create a “Fred Number”, Fr, accessed with \FR, you’d use the command:

\newnondim{\FR}{Fr}

If you would like the nondimensional numbers to appear in a format other than \mathit, simply use \renewcommand to change the \nondimstyle macro to another formatting command. For example, to have non-dimensional numbers appear in roman style, use

\renewcommand{\nondimstyle}{\mathrm}

Let me know if there are other non-dimensional numbers you would like added.

Table 1. Defined Non-dimensional numbers

Number Macro Result
Biot  \BI  Bi
Eckert  \EC  Ec
Fourier  \FO  Fo
Görtler  \GO  Gö
Grashof  \GR  Gr
Nusselt  \NU  Nu
Peclet  \PE  Pe
Prandtl  \PR  Pr
Rayleigh  \RA  Ra
Reynolds \RE  Re
Schmidt  \SC  Sc
Sherwood \SH  Sh
Stanton  \ST  St
Weber  \WE  We

2 Typesetting units

The me-tools package used to provide the \unit and \unitfrac macros for typesetting units. I now recommend the rather nice units package for this instead.

However, for backwards compatibility with older versions of the me-tools package, the meunits package is provided which supplied the \unit and \unitfrac macros.

3 Degree Symbols

The macros \degreesC, \degreesF, and \degreesR are provided for typesetting ^{∘}C, ^{∘}F, and ∘R, respectively. The are defined such that if an existing package has already defined them, the existing definition is not changed.

Table 2. Defined Fraction-like macros

Macro  Results

$\abs{x}$  |x|
$\absfrac\{x\}\{y\}$      |\frac{x}{y}|
$\ddfrac\{T\}\{x\}$       \frac{dT}{dx}
$\ddfrac[2]\{T\}\{x\}$   \frac{d^{2}T}{dx^{2}}
$\ppfrac\{T\}\{z\}$       \frac{∂T}{∂z}
$\ppfrac[2]\{T\}\{z\}$   \frac{∂^{2}T}{∂z^{2}}

4 Fraction Macros The package provides a number of macros for conveniently processing fraction-like constructs. Table 2 shows some examples (note the optional arguments in some of the examples which raise the derivatives to an arbitrary power).

Questions? Recommendations? Mail questions, bugs, and recommendations to kaszeta@me.umn.edu.